OpenStudy (anonymous):

What is the remainder when (x3 − 6x2 − 42x + 79) is divided by (x + 5)?

OpenStudy (welshfella):

use reminder theorem The remainder when P(x) is divided by (x - a) = f(a)

OpenStudy (welshfella):

so here you need to find P(-5)

OpenStudy (anonymous):

i tried plugging in -5, but i kept getting numbers that weren't in the answer choices

OpenStudy (welshfella):

(P(x) is your polynomial)

OpenStudy (welshfella):

oh maybe the choices are all wrong i'll see what i get

OpenStudy (anonymous):

okay thank you

OpenStudy (welshfella):

i get 14

OpenStudy (welshfella):

f(-5) = -125 - 150 - (42*-5) + 79

OpenStudy (anonymous):

thats one of the answer choices! and i was doing +150 idk why lol oops

OpenStudy (anonymous):

can you help me with one more

OpenStudy (welshfella):

ok

OpenStudy (anonymous):

Describe the graph of the function f(x) = x3 − 11x2 + 36x − 36. Include the y-intercept, x-intercepts, and the end behavior.

OpenStudy (welshfella):

well the y-intercept occurs when x = 0 so find f(0) and you get that one

OpenStudy (anonymous):

okay, is the end behavior down?

OpenStudy (welshfella):

hold on

OpenStudy (anonymous):

and i got -36 for the y-intercept

OpenStudy (welshfella):

right

OpenStudy (anonymous):

yay

OpenStudy (welshfella):

its been a while since i did end behaviours fo find the x intercepts we need to solve x^3 -11x^2 + 36x - 36

OpenStudy (anonymous):

So set it equal to 0?

OpenStudy (welshfella):

yes

OpenStudy (welshfella):

as the last number is 36 we could try 2,3 and 6 to see if they are roots

OpenStudy (anonymous):

Can there be more than one answer when you do that tho

OpenStudy (welshfella):

f(2) = 8 - 44 + 72 - 36 = 0 yes x=2 is one x-intercept

OpenStudy (anonymous):

i think they all work

OpenStudy (welshfella):

yes they do thats is pure luck!!!!

OpenStudy (welshfella):

lol

OpenStudy (anonymous):

I think the end behavior is down because it hits x at 3 points

OpenStudy (welshfella):

normally i would divide the polynomial by x - 2 and try to factor the quotient

OpenStudy (welshfella):

no i dont think we can say that as the last zero is 6 try finding f(7) and f(8) and see if they are increasing or decreasing

OpenStudy (welshfella):

f(7) = 20 f(8) = 60 so end behaivour is increasing

OpenStudy (anonymous):

so up? lol

OpenStudy (welshfella):

yes

OpenStudy (anonymous):

yay! thank you so much!

OpenStudy (welshfella):

i'll draw the graph on my calculator to confirm

OpenStudy (welshfella):

yep it looks somthing like